# Mechanism Shift During Post-training from Autoregressive to Masked Diffusion Language Models

### v2

@@ Abstract
Post-training pretrained Autoregressive models (ARMs) into Masked Diffusion models (MDMs) has emerged as a cost-effective strategy to overcome the limitations of sequential generation. However, the internal algorithmic transformations induced by this paradigm shift remain unexplored, leaving it unclear whether post-trained MDMs acquire genuine bidirectional reasoning capabilities or merely repackage autoregressive heuristics. In this work, we address this question by conducting a comparative circuit analysis of ARMs and their MDM counterparts. Our analysis reveals a systematic “mechanism shift” dependent on the structural nature of the task. Structurally, we observe a distinct divergence: while MDMs largely retain autoregressive circuitry for tasks dominated by local causal dependencies, they abandon initialized pathways for global planning tasks, exhibiting distinct rewiring characterized by increased early-layer processing. Semantically, we identify a transition from sharp, localized specialization in ARMs to distributed integration in MDMs. Through these findings, we conclude that diffusion post-training does not merely adapt model parameters but fundamentally reorganizes internal computation to support non-sequential global planning.
@@ 1 Introduction
Large language models have achieved near-human performance across diverse linguistic tasks OpenAI et al. (2024); Qwen et al. (2025); Touvron et al. (2023). Despite these advances, the prevalent autoregressive framework Vaswani et al. (2017); Radford et al. (2019) imposes structural limitations on generation Welleck et al. (2019); Bengio et al. (2015). Specifically, sequential generation under causal masking prevents correcting past tokens Gu et al. (2018); Welleck et al. (2019); Vaswani et al. (2017) so that early errors propagate and amplify throughout the sequence, as the model cannot correct past inaccuracies Bengio et al. (2015); Ranzato et al. (2016). Moreover, many reasoning and planning tasks require global reasoning, where early decisions must account for constraints that apply to the entire sequence Gu et al. (2018); Ye et al. (2025a).
Masked diffusion models (MDMs) have gained increasing interest as a non-autoregressive paradigm, with structural properties well-suited to overcoming these limitations Austin et al. (2021); Sahoo et al. (2024). However, training diffusion-based language models from scratch remains computationally expensive due to slower convergence Gong et al. (2025). To mitigate this cost, recent work proposes post-training pretrained autoregressive models (ARMs) to the diffusion paradigm Gong et al. (2025); Ye et al. (2025b). Models such as Dream Ye et al. (2025b) demonstrate that this strategy can achieve strong performance while requiring only a fraction of the compute needed for training from scratch.
@@ 2.1 Masked Diffusion Models
While training such models from scratch Nie et al. (2025) is computationally expensive, recent work has shown that pretrained ARMs can be effectively post-trained into MDMs Gong et al. (2025). Rather than learning diffusion dynamics from scratch, these approaches initialize from a pretrained ARM. The model is then post-trained to iteratively denoise partially masked inputs, instead of predicting the next token autoregressively. From DiffuLLaMA Gong et al. (2025) to Dream Ye et al. (2025b), this line of work shows that post-training pretrained ARMs to MDM objectives can retain many practical advantages of diffusion—such as parallel decoding, iterative refinement, and bidirectional attention—while substantially reducing training cost. Post-trained MDMs also achieve strong performance on directionality-sensitive tasks. However, while these works establish the effectiveness of post-trained MDMs, they largely focus on performance and efficiency, leaving open the question of how diffusion objectives reshape the underlying computational mechanisms.
@@ 2.2 Mechanistic Interpretability and Circuits
Empirical studies show that many behaviors can be explained by sparse circuits involving only a small fraction of model connections Bhaskar et al. (2024); Wang et al. (2023). Methods such as Edge Attribution Patching identify these subgraphs via gradient-based attribution, enabling circuit discovery for tasks including indirect object identification and numerical comparison Hanna et al. (2023); Lieberum et al. (2023); Bhaskar et al. (2024). Across models and scales, similar circuits—such as induction heads—recur consistently, suggesting that they implement stable algorithmic functions rather than incidental patterns Prakash et al. (2024); Tigges et al. (2024); Ou et al. (2025); Wang et al. (2025). Recent automated approaches, including ACDC and EAP, further enable scalable circuit discovery without manual inspection Syed et al. (2024); Bhaskar et al. (2024).
While mechanistic analyses have begun to probe diffusion models in the vision domain Shi et al. (2025); Niedoba et al. (2025), comparable studies for text diffusion models remain limited. As a result, it is unclear whether diffusion objectives induce distinct computational strategies in language models or primarily reorganize existing autoregressive circuitry.
@@ 3 Method
To investigate the mechanistic shift from ARMs to MDMs, we adopt a comparative framework. Our primary objective is to distinguish the impact of shifting the learning objective from causal modeling to masked diffusion on circuit topology. We control for architectural confounding by analyzing ARMs alongside their directly post-trained MDM counterparts.
@@ 3.1 Models and Configuration
To enable a direct comparison, we pair each standard autoregressive checkpoint with an MDM post-trained from that specific ARM base:
@@ 3.2 Tasks and Datasets
We analyze circuit behavior on two distinct tasks to highlight different performance capabilities: one focused on autoregressive causal dependencies and another on global reasoning. Additional details for tasks and datasets are provided in Appendix A.
@@
A canonical interpretability task primarily solved by Induction Heads in ARMs Wang et al. (2023). We select this task to serve as a representative baseline for causal reasoning; specifically, we aim to investigate whether the induction circuitry—which inherently relies on sequential, left-to-right context to copy tokens—persists, dissolves, or transforms when the training objective shifts from causal prediction to masked diffusion.
@@
A numerical reasoning task where the model is provided with a set of integers and a target value and must generate a valid arithmetic equation using the inputs to equal the target Ye et al. (2025a). This task serves as a critical benchmark for global reasoning because it requires inverse planning—decisions made early in the sequence must be strictly conditioned on the final goal—thereby challenging the left-to-right causality of ARMs while favoring the bidirectional context and global refinement capabilities of MDMs.
Inference Configuration
For all experiments, we use task-dependent generation lengths. For IOI, we use a single diffusion step, while for Countdown, we align diffusion steps with the target sequence length. Details are provided in Appendix A.
@@ 3.3 Circuit Discovery and Analysis Pipeline
We hypothesize that the transition to masked diffusion induces specific algorithmic shifts in tasks where MDMs outperform ARMs (e.g., Countdown). To verify this, we employ a three-stage pipeline: Discovery, Topological Comparison, and Mechanism Interpretation. This workflow allows us to first locate where the computation changes and then analyze what those changes represent semantically.
@@ 3.3.2 Attribution-Guided Circuit Comparison
To localize where algorithmic changes concentrate when transitioning from ARMs to MDMs, we leverage EAP-IG to compare circuits at the level of edges and their incident components. Rather than treating all discovered edges equally, we focus on those that carry the highest attribution mass under EAP-IG and then identify which components are repeatedly used as sources or sinks of these high-attribution edges.
@@
For each model (ARM and MDM), we first select the set of top-attribution edges (1000 edges), and , by ranking edges according to their EAP-IG scores on identical prompts. We then quantify overlap using the Jaccard similarity . High overlap suggests that the MDM reuses the same high-attribution pathways as the ARM, whereas low overlap indicates that diffusion training recruits a distinct set of edges to implement its generative computation. The choice of selecting the top 1000 edges reflects a trade-off between attribution coverage and circuit sparsity and is empirically justified in Appendix B.
@@
To move from edges to components, we assign each node a score by aggregating the EAP-IG scores of all incident edges, defined as the sum of incoming attributions and outgoing attributions , where denotes the set of top-attribution edges for the model. In words, a component is important if it repeatedly appears as either the source or the target of high-scoring EAP-IG edges. We then define the Top-K Components (with ) for each model as the nodes with the largest . The choice of is empirically motivated and discussed in Appendix B.
Based on these metrics, we first assess the degree of circuit reuse between ARMs and MDMs via overlap of high-attribution edges. High edge overlap provides evidence of mechanistic recycling, where the MDM relies on pre-trained autoregressive pathways, while low overlap indicates the emergence of diffusion-specific generative strategies. Within these regions, the Top-K EAP-IG Components pinpoint nodes that sit at the endpoints of the most influential edges, yielding a focused set of components for downstream mechanistic analysis via logit lens and neuron-level visualization.
@@ 3.3.3 Mechanism Interpretation
Having identified where the circuit changes, we investigate how the computation differs by applying interpretability techniques specifically to the Top-K Divergent Components:
@@
Following the logit lens framework nostalgebraist (2020), we project intermediate activations into the vocabulary space via the unembedding matrix , yielding and , where indexes a component (e.g., an attention head), denotes the token position, and the diffusion timestep.
For autoregressive models (Qwen, LLaMA), we apply the component-wise logit lens to all token positions in the sequence on both IOI and Countdown tasks. This enables us to trace how the model’s internal representation becomes aligned with the target output across the sequence and to decompose the contributions of the residual stream, attention heads, and MLPs at each token position.
For MDMs (Dream, DiffuLLaMA), we examine how these component-wise projections evolve across the diffusion time steps at a fixed set of token positions. Dream supports per-head and per-MLP decompositions, while DiffuLLaMA is analyzed at the level of residual, attention, MLP, and residual-out layers.
This analysis allows us to assess whether a given component has shifted its semantic role—for example, from contributing primarily to next-token prediction in an ARM to encoding a global target or distant operator in an MDM—and how such semantic alignment emerges progressively over diffusion time. Importantly, when applying the unembedding matrix to intermediate component activations, we do not interpret the resulting logits as the model’s final predictions. Instead, they serve as a diagnostic probe that reveals which tokens a component is linearly aligned with at a particular stage of computation.
@@
In this work, we define a “neuron” as a single scalar coordinate in the model’s residual stream at a given transformer layer (i.e., one element of the hidden dimension of the layer output). For both the IOI and Countdown settings, we record activations for all combinations of layers, neuron indices, and token positions on the evaluation data. Following the methodology of Bills et al. (2023), we then extract the input tokens that elicit the largest-magnitude activations for a given neuron and use these highly activating examples for qualitative inspection and automated explanation.
We focus this fine-grained analysis on the Top-K Divergent Components identified in the previous step: specifically, for divergent MLP layers, we visualize the top activating neurons to understand the specific attributes they extract; for divergent Attention layers, we analyze the head’s output features to determine what information is being moved. The primary goal of this visualization is to characterize the semantic feature selectivity of the circuit. By mapping high-activation neurons to their corresponding tokens, we aim to determine if MDM components have learned to encode non-causal features (e.g., attending to or encoding “future target” tokens available during the diffusion process) that are fundamentally inaccessible to the standard autoregressive model.
@@ 4.1 Circuit-Level Differences
| Setting | Edge Overlap | Top- Overlap |
| IOI (Qwen / Dream) | 0.193 | 0.105 |
| IOI (LLaMA-2 / DiffuLLaMA) | 0.088 | 0.124 |
| Countdown (Qwen / Dream) | 0.008 | 0.093 |
| Countdown (LLaMA-2 / DiffuLLaMA) | 0.032 | 0.081 |
Figure 2 illustrates the circuit structures extracted for both ARMs and MDMs across the IOI and Countdown tasks, where each circuit represents the average structure aggregated over multiple prompts and diffusion steps.
For the IOI task, the overall circuit topology of Dream closely resembles that of Qwen, with dominant interactions concentrated in similar layers and mediated by comparable attention pathways. A similar pattern is observed for DiffuLLaMA relative to its LLaMA-2 counterpart. This indicates that, for tasks primarily driven by local or causal dependencies—such as IOI, which effectively requires only shallow or single-step generation—post-trained MDMs largely preserve the autoregressive circuitry inherited from their ARM initializations.
This qualitative similarity is further supported by quantitative circuit similarity metrics reported in Table 1. Across both model families, the IOI task exhibits consistently higher overlap in high-attribution edges and Top- components compared to the Countdown task, indicating substantial mechanistic reuse rather than the emergence of new diffusion-specific circuitry in tasks dominated by local dependency structures.
In contrast, the Countdown task exhibits a clear shift in circuit organization between ARMs and MDMs. In the Countdown task, the ARM circuit becomes increasingly dense toward mid-to-late layers, whereas the MDM circuit concentrates a larger fraction of interactions in early layers, resulting in a front-loaded organization of computation. This redistribution is consistently observed both in the aggregated circuit statistics (Figure 2) and in the representative circuit visualizations (Figure 3), indicating a systematic reallocation of computation under the diffusion objective.
This depth-wise shift toward increased early-layer engagement is consistently observed across both model families, with Dream and DiffuLLaMA exhibiting increased early-layer engagement compared to their respective autoregressive baselines. Quantitatively, this divergence is reflected in markedly lower edge-level and component-level overlap scores for the Countdown task (Table 1), indicating that diffusion post-training induces non-trivial reorganization of the underlying circuitry.
Together, these results reveal a clear dichotomy in how masked diffusion post-training reshapes internal circuitry. For tasks dominated by local, causal dependencies such as IOI, post-trained MDMs retain a substantial fraction of the original autoregressive circuitry. High overlap in both edge-level attribution and Top- components indicates that the causal reasoning machinery learned during autoregressive pre-training remains functionally effective and is therefore reused rather than replaced.
In contrast, for tasks requiring global constraint satisfaction, such as Countdown, the autoregressive circuitry proves insufficient. Here, diffusion post-training induces the emergence of distinct early-layer components that are largely absent from the original ARM circuits, accompanied by a marked collapse in circuit overlap. These newly emphasized early-layer components become the primary substrates for computation, suggesting that MDMs restructure their internal mechanisms to support global planning by front-loading computation rather than relying on sequential, causal pathways.
Taken together, these findings demonstrate that masked diffusion post-training does not simply overwrite autoregressive mechanisms. Instead, it selectively preserves causal circuitry where it remains compatible with the task, while inducing genuinely new, diffusion-specific components when global reasoning demands exceed the expressive capacity of autoregressive computation.

### v3

@@ Abstract
Post-training pretrained autoregressive models (ARMs) into masked diffusion models (MDMs) has emerged as a cost-effective way to overcome the limitations of sequential generation. Yet the internal algorithmic changes induced by this shift remain poorly understood, leaving it unclear whether post-trained MDMs acquire genuine bidirectional reasoning or merely repackage autoregressive heuristics. We address this question through a comparative circuit analysis of ARMs and their MDM counterparts. Our analysis reveals a systematic “mechanism shift” that depends on the structural nature of the task. MDMs largely preserve autoregressive circuitry for tasks driven by local causal dependencies, but for global planning tasks they abandon initialized pathways and exhibit distinct rewiring with increased early-layer processing. At the semantic level, we observe a transition from sharp, localized specialization in ARMs to distributed integration in MDMs. These findings show that diffusion post-training does not simply adjust model parameters, but reorganizes internal computation to support non-sequential global planning.
@@ 1 Introduction
Large language models have achieved near-human performance across diverse linguistic tasks OpenAI et al. (2024); Qwen et al. (2025); Touvron et al. (2023). Despite these advances, the prevalent autoregressive framework Radford et al. (2019) imposes structural limitations on generation Welleck et al. (2019); Bengio et al. (2015). Specifically, sequential generation under causal masking prevents correcting past tokens Gu et al. (2018); Welleck et al. (2019); Vaswani et al. (2017) so that early errors propagate and amplify throughout the sequence, as the model cannot correct past inaccuracies Bengio et al. (2015); Ranzato et al. (2016). Moreover, many reasoning and planning tasks require global reasoning, where early decisions must account for constraints that apply to the entire sequence Gu et al. (2018); Ye et al. (2025a).
Masked diffusion models (MDMs) have gained increasing interest as a non-autoregressive paradigm, with structural properties well-suited to overcoming these limitations Austin et al. (2021); Sahoo et al. (2024). However, training diffusion-based language models from scratch remains computationally expensive due to slower convergence Gong et al. (2025). To mitigate this cost, recent work proposes post-training pretrained autoregressive models (ARMs) to the diffusion paradigm Gong et al. (2025); Ye et al. (2025c). Models such as Dream Ye et al. (2025c) demonstrate that this strategy can achieve strong performance while requiring only a fraction of the compute needed for training from scratch.
@@ 2.1 Masked Diffusion Models
While training such models from scratch Nie et al. (2025) is computationally expensive, recent work has shown that pretrained ARMs can be effectively post-trained into MDMs Gong et al. (2025). Rather than learning diffusion dynamics from scratch, these approaches initialize from a pretrained ARM. The model is then post-trained to iteratively denoise partially masked inputs, instead of predicting the next token autoregressively. From DiffuLLaMA Gong et al. (2025) to Dream Ye et al. (2025c), this line of work shows that post-training pretrained ARMs to MDM objectives can retain many practical advantages of diffusion—such as parallel decoding, iterative refinement, and bidirectional attention—while substantially reducing training cost. Post-trained MDMs also achieve strong performance on directionality-sensitive tasks. However, while these works establish the effectiveness of post-trained MDMs, they largely focus on performance and efficiency, leaving open the question of how diffusion objectives reshape the underlying computational mechanisms.
@@ 2.2 Mechanistic Interpretability and Circuits
Empirical studies show that many behaviors can be explained by sparse circuits involving only a small fraction of model connections Bhaskar et al. (2024); Wang et al. (2023). Methods such as Edge Attribution Patching identify these subgraphs via gradient-based attribution, enabling circuit discovery for tasks including indirect object identification and numerical comparison Hanna et al. (2023); Lieberum et al. (2023); Bhaskar et al. (2024). Across models and scales, similar circuits,such as induction heads recur consistently, suggesting that they implement stable algorithmic functions rather than incidental patterns Prakash et al. (2024); Tigges et al. (2024); Ou et al. (2025); Wang et al. (2025). Recent automated approaches, including ACDC and EAP, further enable scalable circuit discovery without manual inspection Syed et al. (2024); Bhaskar et al. (2024).
While mechanistic analyses have begun to probe diffusion models in the vision domain Shi et al. (2025); Niedoba et al. (2025), comparable studies for text diffusion models remain limited. As a result, it is unclear whether diffusion objectives induce distinct computational strategies in language models or reorganize existing autoregressive circuitry.
@@ 3 Method
To investigate the mechanistic shift from ARMs to MDMs, we adopt a comparative framework. We analyze pretrained ARMs alongside their directly post-trained MDM counterparts, allowing us to isolate the effect of changing the learning objective from causal modeling to masked diffusion while controlling for architectural confounding.
@@ 3.1 Models and Configuration
To enable a direct comparison, we pair each autoregressive checkpoint with its corresponding post-trained MDM derived from the same backbone. Concretely, we compare Qwen2.5-7B with Dream-Base-7B Ye et al. (2025c), which is post-trained on the Qwen backbone, and LLaMA-2-7B with DiffuLLaMA-7B Gong et al. (2025), which is initialized from LLaMA 2 weights.
@@ 3.2 Tasks and Datasets
We select two tasks that lie at opposite ends of the local–global reasoning spectrum: one dominated by autoregressive causal dependencies and one requiring global constraint integration. Let denote the target token at position , the prefix context, and the full sequence context excluding . Using Dream-based conditional entropy, we define a task as exhibiting more local reasoning when , and more global reasoning when . Intuitively, the former indicates that prefix context is largely sufficient, whereas the latter indicates that bidirectional full-sequence context substantially reduces uncertainty. Additional details are provided in Appendix A.
@@
A canonical interpretability task primarily solved by Induction Heads in ARMs Wang et al. (2023). We use this task as a representative benchmark for local reasoning, asking whether induction circuitry—which relies on sequential, left-to-right context to copy tokens—is preserved, dissolved, or transformed when the learning objective shifts from causal prediction to masked diffusion.
@@
A numerical reasoning task in which the model is given a set of integers and a target value, and must generate a valid arithmetic expression that equals the target Ye et al. (2025a). We use this task as a representative benchmark for global reasoning, since it requires inverse planning: decisions made early in the sequence must be conditioned on the final goal. This structure challenges the left-to-right causality of ARMs while favoring the bidirectional context and iterative refinement capabilities of MDMs Ye et al. (2025b).
Inference Configuration:
We use task-dependent generation lengths in all experiments. For IOI, we use a single diffusion step; for Countdown, we align the number of diffusion steps with the target sequence length. Additional details are provided in Appendix A.
@@ 3.3 Circuit Discovery and Analysis Pipeline
We hypothesize that the transition to masked diffusion induces specific algorithmic shifts in tasks where MDMs outperform ARMs (e.g., Countdown). To verify this, we employ a three-stage pipeline: Discovery, Topological Comparison, and Mechanism Interpretation.
@@ 3.3.2 Attribution-Guided Circuit Comparison
To localize where algorithmic changes concentrate when transitioning from ARMs to MDMs, we leverage EAP-IG to compare circuits at the level of edges and the components they connect. Rather than treating all discovered edges equally, we focus on those that carry the highest attribution mass under EAP-IG and then identify which components are repeatedly used as sources or sinks of these high-attribution edges.
@@
For each model (ARM and MDM), we first select the set of top-attribution edges (1000 edges), and , by ranking edges according to their EAP-IG scores on identical prompts. We quantify overlap using the Jaccard similarity . High overlap suggests that the MDM reuses the same pathways as the ARM, whereas low overlap indicates that diffusion training recruits a distinct set of edges to implement its generative computation. The choice of selecting the top 1000 edges reflects a trade-off between attribution coverage and circuit sparsity and is empirically justified in Appendix B.
@@
To move from edges to components, we assign each node a score by aggregating the EAP-IG scores of all incident edges, defined as the sum of incoming attributions and outgoing attributions , where denotes the set of top-attribution edges for the model. Thus, a component is considered important if it repeatedly appears as either the source or the target of high-scoring EAP-IG edges. We then define the Top-K Components (with ) for each model as the nodes with the largest . The choice of is empirically motivated and discussed in Appendix B.
Layer-wise Center of Gravity (CoG):
To quantify where the discovered computation is concentrated across depth, we compute a layer-wise Center of Gravity (CoG) from the EAP-IG attribution profile. Let denote the total EAP-IG attribution mass associated with layer in the discovered circuit, and define . Lower CoG indicates more front-loaded computation in earlier layers, while higher CoG indicates greater reliance on middle or late layers. We use CoG as a scalar summary of depth-wise circuit organization. Formal details are provided in Appendix C.
Using these edge-, component-, and layer-level summaries, we first assess the degree of circuit reuse between ARMs and MDMs via overlap of high-attribution edges. High edge overlap provides evidence of mechanistic recycling, where the MDM relies on pre-trained autoregressive pathways, while low overlap indicates the emergence of diffusion-specific generative strategies. Among these edges, Top-K Components pinpoint the nodes at the most influential interactions, yielding a focused set for downstream analysis. CoG then provides a scalar summary of where this computation is depth-wise concentrated, complementing the overlap and component comparisons.
@@ 3.3.3 Mechanism Interpretation
Having identified where the circuit changes, we investigate how the computation differs by applying interpretability techniques specifically to the Top-K EAP-IG Components:
@@
Following the logit lens framework nostalgebraist (2020), we project intermediate activations into vocabulary space via the unembedding matrix . Since each MDM is obtained by post-training from the same pretrained backbone as its paired ARM, we use as a shared linear readout basis across both models: , , where indexes a component, the token position, and the diffusion timestep. We treat these not as the model’s native predictive distributions, but as diagnostic probes of which vocabulary directions each component is aligned with under a shared semantic frame.
For autoregressive models (Qwen, LLaMA), we apply the component-wise logit lens to all token positions in the sequence on both IOI and Countdown tasks. This enables us to trace how the model’s representation becomes aligned with the target output across the sequence and to decompose the contributions of the residual stream, attention heads, and MLPs at each token position. For MDMs (Dream, DiffuLLaMA), we examine how these component-wise projections evolve across the diffusion time steps at a fixed set of token positions. Dream supports per-head and per-MLP decompositions, while DiffuLLaMA is analyzed at the level of residual, attention, MLP, and residual-out layers.
This analysis allows us to assess whether a given component has shifted its semantic role—for example, from contributing primarily to next-token prediction in an ARM to encoding a global target or distant operator in an MDM—and how such semantic alignment emerges progressively over diffusion time. Importantly, when applying the unembedding matrix to intermediate component activations, we do not interpret the resulting logits as the model’s final predictions. Instead, they serve as a diagnostic probe that reveals which tokens a component is linearly aligned with at a stage of computation.
To complement qualitative inspection with scalar summaries of semantic specialization, we compute quantitative probes from the component-wise logit distributions. Specifically, we measure (i) NameFrac@K, the fraction of person-name tokens among the top- aligned tokens, where person-name tokens are identified using a pretrained BERT-based named entity recognition (NER) model filtered for the PERSON label; (ii) LogitGap, the gap between the top-1 and top-2 logits, which captures selective amplification; (iii) the log-mean-exp dominance gap (LME), which compares the dominance of name versus non-name tokens within the top- set; and (iv) top- entropy, which quantifies whether alignment is sharply concentrated or broadly distributed. Together, these probes provide an objective characterization of component-level specialization and prediction dominance. Formal definitions and details are provided in Appendix C.
@@
In this work, we define a “neuron” as a single scalar coordinate in the model’s residual stream at a given transformer layer (i.e., one element of the hidden dimension of the layer output). For both the IOI and Countdown settings, we record activations for all combinations of layers, neuron indices, and token positions on the evaluation data. Following the methodology of Bills et al. (2023), we then extract the input tokens that elicit the largest-magnitude activations for a given neuron and use these activating examples for qualitative inspection and automated explanation.
We focus this fine-grained analysis on the Top-K EAP-IG components identified in the previous step: specifically, for high-attribution MLP layers, we visualize the top activating neurons to understand the specific attributes they extract; for high-attribution Attention layers, we analyze the head’s output features to determine what information is transmitted. The primary goal of this visualization is to characterize the semantic feature selectivity of the circuit. By mapping high-activation neurons to their corresponding tokens, we aim to determine if MDM components have learned to encode non-causal features (e.g., attending to or encoding “future target” tokens available during the diffusion process) that are fundamentally inaccessible to the ARMs.
Together with the component-level logit-lens probes, this neuron-level analysis helps distinguish whether diffusion post-training changes only the distribution of output-aligned predictions or also reshapes the underlying internal features that support those predictions.
@@ 4.1 Circuit-Level Differences
Figure 2 shows the circuits for ARMs and MDMs on IOI and Countdown. Combined with the overlap and CoG metrics in Table 1, the results reveal a task-dependent dichotomy.
For IOI, MDMs largely preserve the autoregressive circuitry inherited from their ARM initializations. Table 1 quantifies this structural retention using both edge overlap and CoG. Although the absolute overlap score appears numerically small (e.g., 0.193 for Qwen/Dream), causal ablation shows that the shared edges form the functional core of the circuit: ablating the intersection reduces IOI accuracy to 28.1%, whereas ablating a random subset of non-shared edges of equal size retains 92.4% ( 1.2%) accuracy across five random seeds. The CoG also remains relatively stable in the mid-to-late layers (17.5 vs. 20.4), indicating that causal reasoning machinery learned during pre-training is reused rather than replaced.
| Task & Models | Overlap | CoG (Layer) | ||
| Edge | Top- | ARM | MDM | |
| IOI Qwen / Dream | 0.193 | 0.105 | 17.5 | 20.4 |
| IOI LLaMA / DiffuLLaMA | 0.088 | 0.124 | 18.2 | 19.8 |
| Countdown Qwen / Dream | 0.008 | 0.093 | 16.5 | 4.8 |
| Countdown LLaMA / DiffuLLaMA | 0.032 | 0.081 | 17.1 | 5.3 |
In contrast, for Countdown, the autoregressive circuitry is no longer sufficient. Circuit overlap drops sharply (e.g., to 0.008 for Qwen/Dream), indicating substantial structural decoupling. This change is accompanied by a strong shift in depth, with the CoG moving from layer 16.5 in the ARM to 4.8 in the MDM. This depth shift is also consistent with the aggregated layer-wise component usage statistics, which show increased concentration in earlier layers (Appendix B Figure 3).
Taken together, these findings show that masked diffusion post-training does not simply overwrite autoregressive mechanisms. Instead, it preserves causal circuitry where it remains task-compatible, while inducing new early-layer diffusion components when global reasoning exceeds the limits of sequential autoregressive computation.